OpenStudy (anonymous):

Let f(x) be a quadratic function. If \(x^2 - 4x + 6 ≤ f(x) ≤ 2x^2 - 8x +10\) for all real numbers x, and f(12) = 182, find f(17).

OpenStudy (turingtest):

my first thought is to try factoring and see if that takes us somewhere

OpenStudy (anonymous):

Aren't they both unfactorable?

OpenStudy (turingtest):

I guess, I'm just brainstorming (and multitasking)

OpenStudy (turingtest):

changing the second into 2(x^2-4x+5) may shed some light is thought #2

OpenStudy (turingtest):

let's see\[f(12)=102\le182\le202\]

OpenStudy (turingtest):

hm...

OpenStudy (turingtest):

I think I need to stop multitasking to do this

OpenStudy (anonymous):

\[227 ≤ f(17) ≤ 462\]

OpenStudy (turingtest):

I have a strong feeling that @asnaseer is going to save us

OpenStudy (asnaseer):

could we make use of the fact that if the LHS=RHS then that fixes f(x)?

OpenStudy (asnaseer):

i.e.\[x^2-4x+6=2x^2-8x+10\]f we solve this we get:\[(x-2)^2=0\implies x=2\]

OpenStudy (asnaseer):

so at x=2 both the LHS (left hand side) and RHS (right hand side) are equal

OpenStudy (turingtest):

interesting...

OpenStudy (anonymous):

That would make f(2) = 2?

OpenStudy (anonymous):

When you plug it back in anyway...

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

and if we define f(x) as:\[f(x)=ax^2+bx+c\]then maybe we can form 3 equations and solve for a, b and c

OpenStudy (asnaseer):

we have two equations so far

OpenStudy (asnaseer):

maybe make use of f(0)?

OpenStudy (asnaseer):

which gives us:\[6\le c\le 10\]

OpenStudy (maheshmeghwal9):

\[x^2 - 4x + 6 ≤ (ax^2+bx+c)≤ 2x^2 - 8x +10\] \[\color{blue}{a(12)^2+b(12)+c}=182=\color{blue}{144a+12b+c.}\] \[\color{red}{a(17)^2+b(17)+c=?}\]\[6 \le c \le 10.\]

OpenStudy (anonymous):

I just looked up how they did it...it's very strange, and I'm having a bit of trouble understanding a part of it...do you want me to post it or would you like to try figuring the rest out?

OpenStudy (asnaseer):

f(1) is also interesting, it gives us:\[3\le a+b+c\le4\]

OpenStudy (asnaseer):

what do you mean by /strange/?

OpenStudy (asnaseer):

is there a particular approach that is used?

OpenStudy (anonymous):

Well, the way they did it is a lot simpler and way different from this way...they completed the square...

OpenStudy (asnaseer):

hmmm... let me think over that approach...

OpenStudy (maheshmeghwal9):

Becoz we all are doing this type of question first time. & they had done it already. I really appreciate @asnaseer 's work:D

OpenStudy (slaaibak):

all three equations have a turning point (minimum) at x=2. so you could use that.

OpenStudy (asnaseer):

I can't /see/ how completing the square helps. but I wonder if we can make use of my last result for f(1) to get 3 possible equations to solve.\[144a+12b+c=182\tag{1}\]\[4a+2b+c=2\tag{2}\]and then either:\[a+b+c=3\tag{3a}\]or:\[a+b+c=4\tag{3b}\]

OpenStudy (anonymous):

Does it become a geometric question when they complete the square?